what is the simplest formula of the compound which has the following percentage composition : carbon 80%, hydrogen 20%, if the molecular mass is 30, calculate the molecular formula.

Grade:11

1 Answers

Viresh kumar
6 years ago
Let there be 100 gram of substance 
Then,C is 80 gram and Hydrogen is 20 Gram
Then,COmpound is:
C80/12H20/1
C6.67H20
Dividing by 6.67
=> CH3(This is Emperical Formula)
Molecular Mass of CH3 is 12 + 3 = 15
M.M. of Compound = 30
So,n=30/15=2
Now multiplying the Emperical Formula by 2,
=> C2H6